OpenStudy (swag):

Please solve this exponent. Please show the work so i can understand (3a^4)^3

OpenStudy (anonymous):

\[(3 \times a^{4})^{3}=3^{3} \times (a^{4})^{3}=3^{3} \times a^{4 \times 3}=27 \times a^{12}\]

OpenStudy (swag):

Thank You

OpenStudy (dmezzullo):

yep thats right

OpenStudy (swag):

LOLOLOLOLOLOLOL I still cant

OpenStudy (swag):

There we go, it fixed

OpenStudy (dmezzullo):

wat lol?